Y varies directly to x^2. If x is decreased
by 200%, find the percentage decrease in y.

My solution-- y=kx^2
1. X itself is 100% so if decrease by 200%,
the new x will be x-x-x= -x
2. The new y = k (-x)^2------= kx^2
3. Percent decrease =. Kx^2- kx^2)/ kx^2
Step 3 made me confused if y= 0%

Answer by Alan3354(69443)

Y varies directly to x^2. If x is decreased
by 200%, find the percentage decrease in y.

My solution-- y=kx^2
1. X itself is 100% so if decrease by 200%,
the new x will be x-x-x= -x
2. The new y = k (-x)^2------= kx^2
3. Percent decrease =. Kx^2- kx^2)/ kx^2
Step 3 made me confused if y= 0%
=================
It is zero. The value of y is the same at x and -x.